The \underline{MIT Lindy Hop Society}, \underline{MIT Ballroom Dance Club} and \underline{MIT Casino Rueda} proudly present

BREAD & BONES

A Fusion Social Dance and Concert

featuring

Amy Kucharik and Friends with Benefits

Saturday, May 13, 2017

Lesson 7:30 p.m. Music and Social Dancing 8:30 p.m.–12:30 a.m.

Walker Memorial, 142 Memorial Dr., Cambridge

$3 MIT students / $5 Other students / $10 General public Includes lesson and complimentary delicious refreshments.

swing@mit.edu \underline{http://mit.edu/swing} Facebook event \underline{http://mit.edu/swing/MKF}

BREAD & BONES

A Fusion Social Dance and Concert

featuring

Amy Kucharik and Friends with Benefits

Saturday, May 13, 2017

Lesson 7:30 p.m. Music and Social Dancing 8:30 p.m.–12:30 a.m.

Walker Memorial, 142 Memorial Dr., Cambridge

$3 MIT students / $5 Other students / $10 General public Includes lesson and complimentary delicious refreshments.

swing@mit.edu \underline{http://mit.edu/swing} Facebook event \underline{http://mit.edu/swing/MKF}

BREAD & BONES

A Fusion Social Dance and Concert

featuring

Amy Kucharik and Friends with Benefits

Saturday, May 13, 2017

Lesson 7:30 p.m. Music and Social Dancing 8:30 p.m.–12:30 a.m.

Walker Memorial, 142 Memorial Dr., Cambridge

$3 MIT students / $5 Other students / $10 General public Includes lesson and complimentary delicious refreshments.

swing@mit.edu \underline{http://mit.edu/swing} Facebook event \underline{http://mit.edu/swing/MKF}

BREAD & BONES

A Fusion Social Dance and Concert

featuring

Amy Kucharik and Friends with Benefits

Saturday, May 13, 2017

Lesson 7:30 p.m. Music and Social Dancing 8:30 p.m.–12:30 a.m.

Walker Memorial, 142 Memorial Dr., Cambridge

$3 MIT students / $5 Other students / $10 General public Includes lesson and complimentary delicious refreshments.

swing@mit.edu \underline{http://mit.edu/swing} Facebook event \underline{http://mit.edu/swing/MKF}